OpenStudy (anonymous):

[6.03] What is the value of the y variable in the solution to the following system of equations? 5x + 3y = 7 3x - 5y = -23 -1 1 -4 4

OpenStudy (jdoe0001):

if you were to multiply horizontally the equation atop by say "3" and you were to multiply horizontally the equation below by say "-5" what would all that turn into? that is (5x + 3y = 7 ) \(\large \times 3\) (3x - 5y = -23) \(\large \times -5\)